OpenStudy (anonymous):

In order to qualify for a girls' regional 1500 meter race, an athlete's personal best time for the season must be under the qualifying time of 5 minutes 42 second. 1. Write an equation that expresses 'd' as the difference of the athlete's personal best time 't' (in seconds) and the qualifying time (in seconds) 2.Make a table that gives the value of 'd' for 't' = 341.7, 343.8, 340.9 , and 342.7. Which make values of 't' in the table are under the qualifying time? How can you tell from the differences?

OpenStudy (anonymous):

5'42'' = 5*60 +42 = 342 seconds 1/ d = t - 342 2/ t | 341.7 | 343.8 | 340.9 -------|---------------|-------------|----------- d | -0.3 | 1.8 | -1.1 The negative values of d make t under the qualifying time.